OpenStudy (anonymous):

PLEASE HELP! LAST HOMEWORK QUESTION! WILL GIVE MEDAL! (On geometric sequences) The question is in the comments, any help is appreciated but if you could please show the working out, that would be great! Thanks!

OpenStudy (anonymous):

a sequence is defined by \[tn =4\times \left( \frac{ 1 }{ 2 } \right)^{n-1}\] a) prove that the sequence is geometric b) find a and r c) find the 7th term as a decimal number